OpenStudy (anonymous):

in what ratio does the plane {2x+2y-2z=1} divide the line segment joining (2,1,5) and (3,4,3)

OpenStudy (anonymous):

let the plane divide the line in the ratio k:1 let (x,y,z) be its co-ordinates,then (x,y,z)=((3k+2)/(k+1),(4k+1)/(k+1),(3k+5)/(k+1)) respectively if this point lies on the plane 2x+2y-2z=1,then 2(3k+2)/(K+1)+2(4k+1)/(k+1)-2(3k+5)/(k+1)=1 6k+4+8k+2-6k-10=k+1 6k+8k-6k-k=1-4-2+10 7k=6 k=6/7 Hence ratio is 6/7:1 or6:7

OpenStudy (anonymous):

Thanks a lot dude =D i previously tried the same thing and got a rediculous answer :P turns out i made an addition mistake :P lol

OpenStudy (anonymous):

yw